I am having trouble understanding this problem. Can you please walk me through the steps using the BCA and ICE tables? Calculate the pH of each of the solutions and the change in pH to 0.01 pH units caused by adding 10.0 mL of 3.68-M HCl to 300. mL of each of the following solutions Change is defined as final minus initial, so if the pH drops upon mixing the change is negative a) water pH before mixing7.00 pH after mixing 0.93 pH change- 6.07 b) 0.158 M C2H3021 pH before mixing 8.97 pH after mixing4.48 pH change- c) 0.158 M HC2H302 pH before mixing 2.77 pH after mixing0.93 pH change - 1.84 d) a buffer solution that is 0.158 M in each C2H32 and HC2H302 pH before mixing- pH after mixing pH change -
Answer #1

b) Consider the ionization of C2H3O21- as below.

C2H3O21- (aq) + H2O (l) --------> HC2H3O2 (aq) + OH- (aq)

Since OH- is formed, we work with Ka of C2H3O21-. We know that Ka (HC2H3O2) = 1.8*10-5. We further know that

Ka*Kb = Kw

where Kb is the base ionization constant of C2H3O21- and Kw = 1.0*10-14 is a constant. Therefore,

Kb = Kw/Ka = (1.0*10-14)/(1.8*10-5)

= 5.55*10-10

Next, write down the expression for Kb as

Kb = [HC2H3O2][OH-]/[C2H3O21-] = (x).(x)/(0.158 – x)

Since Kb is small, we can assume x << 0.158 M and thus,

Kb = 5.55*10-10 = x2/(0.158)

====> x2 = 5.55*10-10*0.158 = 8.769*10-11

====> x = 9.364*10-6

Therefore, [OH-] = 9.364*10-6 M and pOH = -log [OH-] = -log (9.364*10-6 M) = 5.028.

We know that pH + pOH = 14; therefore,

pH = 14 – pOH = 14 – 5.028 = 8.972 ? 8.97.

This denotes the initial pH of the solution.

Now, millimoles C2H3O21- = (volume of C2H3O21- in mL)*(molarity of C2H3O21-) = (300. mL)*(0.158 M) = 47.4 mmole.

Millimoles of HCl added = (volume of HCl in mL)*(molarity of HCl) = (10.0 mL)*(3.68 M) = 36.8 mmole.

C2H3O21- reacts with HCl as per the equation

C2H3O21- (aq) + HCl (aq) --------> HC2H3O2 (aq) + Cl- (aq)

Cl- is the counter ion. C2H3O21- and HC2H3O2 form a buffer.

As per the stoichiometry of the reaction,

millimoles C2H3O21- neutralized = millimoles HCl added = millimoles HC2H3O2 formed = 36.8 mmole.

Millimoles of C2H3O21- retained at equilibrium = (47.4 – 36.8) mmole = 10.6 mmole.

Since the total volume of the solution remains stay, we can express the ratio of the concentrations of C2H3O21- and HC2H3O2 in terms of the number of moles. We determine the pH using the Henderson-Hasslebach equation as

pH = pKa + log [C2H3O21-]/[HC2H3O2]

====> pH = -log (Ka) + log (moles C2H3O21- at equilibrium)/(moles HC2H3O2 at equilibrium)

====> pH = -log (1.8*10-5) + log (10.6 mmole)/(36.8 mmole)

====> pH = 4.745 + log (0.288)

====> pH = 4.745 + (-0.541)

====> pH = 4.204 ? 4.20

This denotes the pH of the solution after the addition of HCl. Therefore, pH change = (final pH) – (initial pH) = (4.20) – (8.97) = -4.77 (ans).

d) We determine the pH using the Henderson-Hasslebach equation as

pH = pKa + log [C2H3O21-]/[HC2H3O2]

====> pH = -log (1.8*10-5) + log (0.158 M)/(0.158 M)

====> pH = 4.745 + log (1)

====> pH = 4.745 + 0 ? 4.74

This denotes the pH of the solution before the addition of HCl.

Now, millimoles C2H3O21- = millimoles of HC2H3O2 = (volume of C2H3O21-/HC2H3O2 in mL)*(molarity of C2H3O21-/HC2H3O2) = (300. mL)*(0.158 M) = 47.4 mmole.

Millimoles of HCl added = (volume of HCl in mL)*(molarity of HCl) = (10.0 mL)*(3.68 M) = 36.8 mmole.

C2H3O21- reacts with HCl as per the equation

C2H3O21- (aq) + HCl (aq) --------> HC2H3O2 (aq) + Cl- (aq)

Cl- is the counter ion. C2H3O21- and HC2H3O2 form a buffer.

As per the stoichiometry of the reaction,

millimoles C2H3O21- neutralized = millimoles HCl added = millimoles HC2H3O2 formed = 36.8 mmole.

Millimoles of C2H3O21- retained at equilibrium = (47.4 – 36.8) mmole = 10.6 mmole.

Millimoles of HC2H3O2 at equilibrium = (47.4 + 36.8) mmole = 84.2 mmole.

Since the total volume of the solution remains stay, we can express the ratio of the concentrations of C2H3O21- and HC2H3O2 in terms of the number of moles. We determine the pH using the Henderson-Hasslebach equation as

pH = pKa + log [C2H3O21-]/[HC2H3O2]

====> pH = 4.745 + log (moles C2H3O21- at equilibrium)/(moles HC2H3O2 at equilibrium)

====> pH = 4.745 + log (10.6 mmole)/(84.2 mmole)

====> pH = 4.745 + log (0.126)

====> pH = 4.745 + (-0.8996)

====> pH = 3.8454 ? 3.84

This denotes the pH of the solution after the addition of HCl. Therefore, pH change = (final pH) – (initial pH) = (3.84) – (4.74) = -0.90 (ans).